OpenStudy (anonymous):

find the missing coordinate? m= -2 (5,3) & (r, -7)+

OpenStudy (raden):

(3-(-7))/(5-r) = -2 (3+7)/(5-r) = -2 10/(5-r) = -2 10 = -2(5-r) 10 = -10 + 2r 10+10=2r 20=2r r=10